# EGG-SR: Embedding Symbolic Equivalence into Symbolic Regression via Equality Graph

### v1

@@ Abstract
Symbolic regression seeks to uncover physical laws from experimental data by searching for closed-form expressions, which is an important task in AI-driven scientific discovery. Yet the exponential growth of the search space of expression renders the task computationally challenging. A promising yet underexplored direction for reducing the effective search space and accelerating training lies in symbolic equivalence: many expressions, although syntactically different, define the same function – for example, $\log(x_{1}^{2}x_{2}^{3})$, $\log(x_{1}^{2})+\log(x_{2}^{3})$, and $2\log(x_{1})+3\log(x_{2})$. Existing algorithms treat such variants as distinct outputs, leading to redundant exploration and slow learning. We introduce EGG-SR, a unified framework that integrates equality graphs (e-graphs) into diverse symbolic regression algorithms, including Monte Carlo Tree Search (MCTS), deep reinforcement learning (DRL), and large language models (LLMs). EGG-SR compactly represents equivalent expressions through the proposed EGG module, enabling more efficient learning by: (1) pruning redundant subtree exploration in EGG-MCTS, (2) aggregating rewards across equivalence classes in EGG-DRL, and (3) enriching feedback prompts in EGG-LLM. Under mild assumptions, we show that embedding e-graphs tightens the regret bound of MCTS and reduces the variance of the DRL gradient estimator. Empirically, EGG-SR consistently enhances multiple baselines across challenging benchmarks, discovering equations with lower normalized mean squared error than state-of-the-art methods. Code implementation is available at: https://www.github.com/jiangnanhugo/egg-sr.
@@ 1 Introduction
Symbolic regression aims to automatically discover physical laws from experimental data and has been widely used in scientific domains (doi:10.1126/science.1165893; journal/2020/aifrynman; cory2024evolving; doi:10.1073/pnas.2413441122). Many contemporary algorithms for symbolic regression formulate the search for optimal expressions as a sequential decision-making process. In literature, existing models learn to predict the optimal sequence of grammar rules (DBLP:conf/iclr/Sun0W023), the pre-order traversal sequence for expression trees (DBLP:conf/iclr/PetersenLMSKK21; DBLP:conf/nips/KamiennydLC22), or directly executable strings that follow Python syntax (DBLP:journals/iclr/llmsr25; DBLP:conf/iclr/ZhangCXB025). This task remains computationally challenging due to its NP-hard nature (journal/tmlr/virgolin2022), that is, the search space of candidate expressions grows exponentially with the data dimension.
@@
A promising yet underexplored direction for reducing the effective search space and accelerating discovery is the integration of symbolic equivalence into learning algorithms. For example, these expressions $\log(x_{1}^{2}x_{2}^{3})$, $\log(x_{1}^{2})+\log(x_{2}^{3})$, and $2\log(x_{1})+3\log(x_{2})$ all represent the same math function and are therefore symbolically equivalent. Ideally, a well-trained model would recognize such equivalences and assign identical probabilities, rewards, or losses to the corresponding predicted expressions (DBLP:conf/icml/AllamanisCKS17), since these expressions produce identical functional outputs and attain the same prediction error on the dataset. In the literature, existing algorithms treat these expressions as distinct outputs, leading to redundant exploration of the search space and slow convergence. The main challenge of this direction is: how to represent equivalent expressions and embed them into modern learning models in a unified and scalable manner?
@@
Since the number of equivalent variants grows exponentially with expression length, explicitly representing the full set of equivalent expressions quickly becomes both time-consuming and memory-intensive. To address this challenge, a line of recent works introduced the equality graph (e-graph), a data structure that compactly encodes the set of equivalent variants by storing shared sub-expressions only once (DBLP:journals/pacmpl/NandiWZWSASGT21; DBLP:journals/pacmpl/WillseyNWFTP21; DBLP:journals/pacmpl/KurashigeJGBNIJ24). E-graphs have since been successfully applied to diverse tasks, including program optimization (DBLP:journals/corr/abs-2407-12794), dataset generation of equivalent expressions (DBLP:conf/naacl/ZhengWGK25), duplicate detection (DBLP:conf/gecco/olivetti25), and expression simplification (DBLP:conf/gecco/FrancaK23). Despite these advances, we find that a general framework that enables diverse symbolic regression algorithms to interact with e-graphs for accelerating the discovery remains missing.
@@
We present a unified framework, Egg-SR, that integrates symbolic equivalence into learning algorithms using e-graphs, to accelerate a broad range of symbolic regression methods. Our framework encompasses Egg-enhanced Monte Carlo Tree Search (Egg-MCTS), Egg-enhanced Deep Reinforcement Learning (Egg-DRL), and Egg-enhanced Large Language model (Egg-LLM). The core idea is to leverage e-graphs to efficiently sample multiple equivalent variants of expressions and compute a new equivalence-aware learning objective. Specifically, (1) in search tree–based methods, Egg-MCTS prunes redundant exploration over equivalent subtrees; (2) in reward-driven learning, Egg-DRL aggregates rewards over equivalent expressions, stabilizing training; (3) in LLM-based approaches, Egg-LLM enriches the feedback prompt with multiple equivalent expressions to better guide next round generation. Under mild theoretical assumptions, we show the benefit of embedding symbolic equivalence into learning: (1) Egg-MCTS offers a tighter regret bound than standard MCTS (DBLP:conf/iclr/Sun0W023), and (2) The gradient estimator of Egg-DRL exhibits a lower variance than that of standard DRL (DBLP:conf/iclr/PetersenLMSKK21).
@@
In experiments, we evaluate Egg-SR with a wide class of symbolic regression baselines across several challenging benchmarks. We demonstrate its advantages over existing approaches using Egg than without. The e-graph module consistently improves performance across diverse frameworks, including MCTS, DRL, and LLM. Egg helps to discover symbolic expressions with lower normalized mean squared error than baseline methods.
@@ 2 Preliminaries
Symbolic Expression. Let $\mathbf{x}=(x_{1},\ldots,x_{n})$ denote input variables and $\mathbf{c}=(c_{1},\ldots,c_{m})$ be coefficients. A symbolic expression $\phi$ connects these variables and coefficients using mathematical operators such as addition, multiplication, and logarithms. For example, $\phi=3\log x_{1}+2\log x_{2}$ is a symbolic expression composed of variables $\{x_{1},x_{2}\}$, operators $\{+,\log\}$, and coefficients $\{c_{1}=3,c_{2}=2\}$. Symbolic expressions can be represented as binary trees (DBLP:conf/gecco/olivetti25), pre-order traversal sequences of the tree (DBLP:conf/iclr/PetersenLMSKK21), or sequences of production rules defined by a context-free grammar (DBLP:conf/iclr/Sun0W023).
To uniformly handle all symbolic objects in this work, we use a context-free grammar to represent symbolic expressions. The grammar is defined by the tuple $\langle V,\Sigma,R,S\rangle$ where (1) $V$ is a set of non-terminal symbols representing arbitrary sub-expressions, such as $V=\{A\}$; (2) $\Sigma$ is a set of terminal symbols, including input variables and coefficients, i.e., $\{x_{1},\dots,x_{n}\}\cup\{\mathtt{const}\}$; (3) $R$ is a set of production rules representing mathematical operations, such as addition or multiplication. For example, $A\to A\times A$ denotes multiplication, and the semantics is to replace the left-hand side with the right-hand side; (4) $S$ is the start symbol, typically set to $S=A$.
By applying a sequence of production rules from the start symbol, each rule sequentially replaces the first non-terminal symbol in the expression. The resulting expression, free of non-terminals, corresponds to a valid expression. Consider an example rule sequence $(A\to A\times A,A\to\texttt{const},A\to\log(x_{1}))$. The first multiplication rule $A\to A\times A$ expands the start symbol $\phi=A$ to $\phi=A\times A$. Applying the second rule $A\to\texttt{const}$ yields $\phi=c_{1}\times A$. Finally, we obtain $\phi=c_{1}\log(x_{1})$.
Symbolic Equivalence under a Rewrite System. Rewrite rules are employed to simplify, rearrange, and reformulate expressions in tasks, such as code optimization and automated theorem proving (huet1980equations; DBLP:journals/pacmpl/NandiWZWSASGT21). A rewrite system provides a systematic procedure for transforming expressions by replacing sub-expressions according to predefined patterns.
Formally, a rewrite rule $r_{i}$ is written as $\mathtt{LHS}\leadsto\mathtt{RHS}$”, where the left-hand side ($\mathtt{LHS}$) specifies a match pattern and the right-hand side ($\mathtt{RHS}$) specifies its replacement. Given a set of rules $\mathcal{R}=\{r_{1},\ldots,r_{L}\}$, the symbolic equivalence relation $\equiv_{\mathcal{R}}$ induced by $\mathcal{R}$ is defined as follows: for two symbolic expressions $\phi_{1}$ and $\phi_{2}$,
@@
where $\phi_{1}\to^{*}\phi_{2}$ denotes that $\phi_{1}$ can be transformed into $\phi_{2}$ through a finite sequence of applications of rules in $\mathcal{R}$. In other words, two expressions are equivalent under $\mathcal{R}$ if one can be rewritten into the other by repeated application of these rules. In this work, we systematically encode known mathematical identities as rewrite rules $\mathcal{R}$ to generate equivalent variants of symbolic expressions.
Figure 1: Applying the rewrite rule $\log(a\times b)\leadsto\log(a)+\log(b)$ to an e-graph representing expression $\log(x_{1}^{3}x_{2}^{2})$. (a) The initialized e-graph consists of e-classes (dashed boxes), each containing equivalent e-nodes (solid boxes). Edges connect e-nodes to their child e-classes. (b) The matching step identifies the e-nodes that match the $\mathtt{LHS}$ of the rule. (c) The substitution step adds new e-classes and edges corresponding to the $\mathtt{RHS}$ to the e-graph. (d) The merging step consolidates equivalent e-classes. The final e-graph in (d) compactly represents two equivalent expressions.
For example, consider the rewrite rule $r_{1}=\log(a\times b)\leadsto\log(a)+\log(b)$, where $a$ and $b$ are placeholders for arbitrary sub-expressions. In our grammar-based representation, $\mathtt{LHS}$ corresponds to the rule sequence $(A\to\log(A),\ A\to A\times A,\ A\to a,\ A\to b)$, and $\mathtt{RHS}$ corresponds to $(A\to A+A,\ A\to\log(A),\ A\to a,\ A\to\log(A),\ A\to b)$. Applying $r_{1}$ to $\phi_{1}=\log(x_{1}^{3}x_{2}^{2})$ yields $\phi_{2}=\log(x_{1}^{3})+\log(x_{2}^{2})$. Thus $\phi_{1}$ and $\phi_{2}$ are symbolically equivalent under $r_{1}$.
Symbolic Regression searches for a closed-form expression from the experimental data, which are widely applied in diverse scientific domains (doi:10.1126/sciadv.abq0279). Given a dataset ${D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, the task searches for the optimal expression $\phi^{*}$, that minimizes the averaged loss on the dataset:
$\phi^{*}\leftarrow\arg\min_{\phi\in\Phi}\;\frac{1}{N}\sum\nolimits_{i=1}^{N}\ell(\phi(\mathbf{x}_{i},\mathbf{c}),y_{i}),$
where $\Phi$ denotes the search space containing all expressions; the loss function $\ell$ measures the distance between the output of the candidate expression $\phi(\mathbf{x}_{i},\mathbf{c})$ and the ground truth $y_{i}$. The coefficient $\mathbf{c}$ within expression $\phi$ are optimized on training data D{D} using optimizers, like BFGS (fletcher2000practical). Since the search space of expressions $\Phi$ is exponentially large in the data dimension, finding the optimal expression is challenging. In fact, this task has been proven NP-hard (journal/tmlr/virgolin2022). Recent works to mitigate this challenge are reviewed in section 4.
@@ 3.1 Egg: Equality graph for Grammar-based Symbolic Expression
Enumerating all equivalent variants of a symbolic expression is combinatorially expensive. Storing these variants explicitly (e.g., in an array) is highly space-inefficient. To mitigate this challenge, we adopt the recently proposed Equality graph (i.e, e-graph) data structure (DBLP:journals/pacmpl/WillseyNWFTP21; DBLP:journals/jar/WaldmannTRB22), which compactly represents the set of equivalent expressions by sharing common subexpressions. We extend e-graphs to support grammar-based symbolic expressions, noted as Egg, facilitating seamless integration with symbolic regression algorithms.
@@
Definition. An e-graph consists of a collection of equivalence classes, called e-classes. Each e-class contains a set of e-nodes representing symbolically equivalent sub-expressions (DBLP:journals/pacmpl/WillseyNWFTP21). Each e-node encodes a mathematical operation and references a list of child e-classes corresponding to its operands. Edges always point from an e-node to e-classes.
@@
Figure 1(d) shows an example e-graph. The color-highlighted part is an e-class (in dashed box) containing two e-nodes (in solid boxes): $A\to\log(A)$” and $A\to A+A$”. The two e-nodes represent logarithmic operation and addition, respectively. The e-node $A\to\log(A)$” has a single edge to its child e-class $A\to A\times A$”, because $\log$-operator involves one operand.
@@
Construction. An e-graph is initialized with a sequence of production rules, representing an input expression. A predefined set of mathematical laws is formatted as rewrite rules (as defined in section 2). Each rule is applied by matching its left-hand side ($\mathtt{LHS}$) pattern against the current e-graph, which involves traversing all e-classes to identify subexpressions that match $\mathtt{LHS}$. For every successful match, new e-classes and e-nodes corresponding to the right-hand side ($\mathtt{RHS}$) of the rule are created. A merge operation is applied to incorporate the new e-class into the matched e-class, thereby preserving the structure of known equivalences. This process, known as equality saturation, iteratively applies pattern matching and merging until either no further rules can be applied or a maximum number of iterations is reached.
@@
Figure 1 shows an example of e-graph construction with the rewrite rule $\log(a\times b)\leadsto\log(a)+\log(b)$, where $a$ and $b$ are placeholders for arbitrary sub-expressions. The e-graph in initialized with an expression $\phi=\log(x_{1}^{3}x_{2}^{2})$ in Figure 1(a). The $\mathtt{LHS}$ is matched against the color-highlighted e-classes in Figure 1(b), where $a=x_{1}^{3}$ and $b=x_{2}^{2}$. The substitute step first constructs e-classes and e-nodes that represent $\mathtt{RHS}$, which are color-highlighted in Figure 1(c). One e-class is merged with the matched e-class in Figure 1(d). The resulting e-graph represents the set of two equivalent expressions: $\log(x_{1}^{3}x_{2}^{2})$ and $\log(x_{1}^{3})+\log(x_{2}^{2})$. The e-graph in Figure 1(d) is space-efficient, because two sub-expressions $x_{1}^{3}$ and $x_{2}^{2}$ are stored only once in the e-graph. Additional Egg visualizations on more complex expressions—including those involving trigonometric and partial-derivative rewrite rules—are provided in Appendix F.1.
@@
Extraction. After the e-graph is saturated, an extraction step is performed to obtain $K$ representative expressions (DBLP:journals/pacmpl/GoharshadyLP24). Because an e-graph encodes up to an exponential number of equivalent expressions, exhaustive enumeration is computationally infeasible. We therefore adopt two practical strategies: (1) cost-based extraction, which selects several simplified expressions by minimizing a user-defined cost function over operators and variables, and (2) random-walk sampling, which generates a batch of expressions by stochastically traversing the e-graph. The detailed procedure is provided in Appendix A.4.
Implementation. Most existing e-graph implementations are either in Haskell (DBLP:journals/pacmpl/WillseyNWFTP21) or provided as Python wrappers (shanabrook2024egglogpythonpythoniclibrary), neither of which is well-suited for symbolic regression. We therefore implement Egg in Python, following the algorithmic pipeline in (DBLP:journals/pacmpl/WillseyNWFTP21). Our implementation supports grammar-based expressions and integrates seamlessly with the sequential decision-making process, which is essential for integration with modern learning algorithms. Egg incorporates a wide class of rewrite rules, including those for logarithmic, trigonometric, and differential operators. Implementation details are provided in Appendix A.
3.2 Egg-SR: Embed Symbolic Equivalence into Symbolic Regression via Egg

### v2

@@ Abstract
Symbolic regression seeks to uncover physical laws from experimental data by searching for closed-form expressions, which is an important task in AI-driven scientific discovery. Yet the exponential growth of the search space of expression renders the task computationally challenging. A promising yet underexplored direction for reducing the search space and accelerating training lies in symbolic equivalence: many expressions, although syntactically different, define the same function – for example, $\log(x_{1}^{2}x_{2}^{3})$, $\log(x_{1}^{2})+\log(x_{2}^{3})$, and $2\log(x_{1})+3\log(x_{2})$. Existing algorithms treat such variants as distinct outputs, leading to redundant exploration and slow learning. We introduce Egg-SR, a unified framework that integrates symbolic equivalence into a class of modern symbolic regression methods, including Monte Carlo Tree Search (MCTS), Deep Reinforcement Learning (DRL), and Large Language Models (LLMs). Egg-SR compactly represents equivalent expressions through the proposed Egg module (via equality graphs), accelerating learning by: (1) pruning redundant subtree exploration in Egg-MCTS, (2) aggregating rewards across equivalent generated sequences in Egg-DRL, and (3) enriching feedback prompts in Egg-LLM. Theoretically, we show the benefit of embedding Egg into learning: it tightens the regret bound of MCTS and reduces the variance of the DRL gradient estimator. Empirically, Egg-SR consistently enhances a class of symbolic regression models across several benchmarks, discovering more accurate expressions within the same time limit.
Project page is at: https://nan-jiang-group.github.io/egg-sr.
@@ 1 Introduction
Symbolic regression aims to automatically discover physical knowledge from experimental data and has been widely used in scientific domains [doi:10.1126/science.1165893, journal/2020/aifrynman, cory2024evolving, doi:10.1073/pnas.2311808121, doi:10.1073/pnas.2413441122]. Many contemporary methods for symbolic regression formulate the search for optimal expressions as a sequential decision-making process. In literature, existing methods learn to predict the sequence of grammar rules [DBLP:conf/iclr/Sun0W023], the traversal sequence of expression trees [DBLP:conf/iclr/PetersenLMSKK21, DBLP:conf/nips/KamiennydLC22], or executable strings that follow Python syntax [DBLP:journals/iclr/llmsr25, DBLP:conf/iclr/ZhangCXB025]. This task remains computationally challenging due to its NP-hard nature [journal/tmlr/virgolin2022], that is, the search space of candidate expressions grows exponentially with the data dimension.
@@
A promising yet underexplored direction for reducing the search space and accelerating discovery is the integration of symbolic equivalence into learning algorithms. For example, these expressions $\log(x_{1}^{2}x_{2}^{3})$, $\log(x_{1}^{2})+\log(x_{2}^{3})$, and $2\log(x_{1})+3\log(x_{2})$ all represent the same math function and are therefore symbolically equivalent. Ideally, a well-trained model would recognize such equivalence and assign identical goodness-of-fit, rewards, or losses to the corresponding predicted expressions [DBLP:conf/icml/AllamanisCKS17], since these expressions produce identical functional outputs and attain the same prediction error on the dataset. In the literature, existing SR algorithms treat these expressions as distinct outputs, leading to redundant exploration of the search space and slow training. The main challenge of this direction is: how to represent symbolically-equivalent expressions and embed them into modern learning frameworks in a unified and scalable manner?
@@
Since the number of equivalent variants grows exponentially with expression length, explicitly maintaining the set of equivalent expressions is not time and memory scalable. To mitigate this scalability challenge, a line of recent works introduced the Equality graph (e-graph), a data structure that compactly encodes the set of equivalent variants by storing shared sub-expressions only once [DBLP:journals/pacmpl/NandiWZWSASGT21, DBLP:journals/pacmpl/WillseyNWFTP21, DBLP:journals/pacmpl/KurashigeJGBNIJ24]. e-graphs have been applied to tasks, including program optimization [DBLP:journals/corr/abs-2407-12794], dataset generation of equivalent expressions [DBLP:conf/naacl/ZhengWGK25]. In genetic programming-based symbolic regression, e-graphs have been used for duplicate detection [DBLP:conf/gecco/olivetti25], expression simplification [DBLP:conf/gecco/FrancaK23], and expression template pattern matching [rEGGression]. Despite the empirical successes, we find that a unified framework that enables diverse symbolic regression algorithms to interact with e-graphs to accelerate learning has room for improvement.
@@
We present a unified framework, Egg-SR, that integrates symbolic equivalence into a class of learning algorithms via e-graphs. Our framework encompasses Egg-MCTS, Egg-DRL, and Egg-LLM. The core idea is to leverage Egg to efficiently sample equivalent variants of expressions predicted by SR algorithms and compute a new equivalence-aware learning objective. Specifically, (1) Egg-MCTS prunes redundant exploration over equivalent subtrees; (2) Egg-DRL aggregates rewards over equivalent expressions, stabilizing training; (3) Egg-LLM enriches the feedback prompt with multiple equivalent expressions to better guide next round predictions. Under mild theoretical assumptions, we show the benefit of embedding symbolic equivalence into learning: (1) Egg-MCTS offers a tighter regret bound than standard MCTS [DBLP:conf/iclr/Sun0W023], and (2) the gradient estimator of Egg-DRL exhibits a lower variance than that of standard DRL [DBLP:conf/iclr/PetersenLMSKK21].
@@
In experiments, we evaluate Egg-SR with several representative symbolic regression baselines across several challenging benchmarks. We demonstrate its advantages over existing approaches using Egg than without. Egg consistently improves performance across diverse frameworks, discovering more accurate expressions than baseline within a fixed time budget.
@@ 2 Preliminaries
Symbolic Expression. Let $\mathbf{x}=(x_{1},\ldots,x_{n})$ denote input variables and $\mathbf{c}=(c_{1},\ldots,c_{m})$ be coefficients. A symbolic expression $\phi$ connects these variables and coefficients using mathematical operators such as addition, multiplication, and logarithm. For example, $\phi=3\log x_{1}+2\log x_{2}$ is a symbolic expression composed of variables $\{x_{1},x_{2}\}$, operators $\{+,\log\}$, and coefficients $\{c_{1}=3,c_{2}=2\}$. In literature, symbolic expressions have been represented as binary trees [DBLP:conf/gecco/olivetti25], pre-order traversal sequences of the binary tree [DBLP:conf/iclr/PetersenLMSKK21], topological traversal sequences of the expression graph [ijcai2024p471, xiang2025graphbased], or sequences of production rules defined by a context-free grammar [DBLP:conf/iclr/Sun0W023].
To handle all symbolic objects in this work, a context-free grammar is adopted to represent symbolic expressions [DBLP:journals/kbs/BrenceTD21]. The grammar is defined by a tuple $\langle V,\Sigma,R,S\rangle$ where (1) $V$ is a set of non-terminal symbols representing arbitrary sub-expressions, i.e., $V=\{A\}$; (2) $\Sigma$ is a set of terminal symbols, including input variables and coefficients, i.e., $\{x_{1},\dots,x_{n}\}\cup\{\mathtt{c}\}$; (3) $R$ is a set of production rules representing mathematical operations. For example, $A\to A\times A$ denotes multiplication, and the semantics is to replace the left-hand side with the right-hand side; (4) $S$ is the start symbol, typically set to $S=A$. Given a sequence of production rules, an expression is constructed by sequentially applying each rule to the leftmost nonterminal, starting from $S$. If the resulting string contains no nonterminals, it corresponds to a valid expression.
Figure 6 (in appendix) shows the expression construction with sequence $(A\to A\times A,A\to\mathtt{c},A\to\log(x_{1}))$. The first rule $A\to A\times A$ expands the start symbol $\phi=A$ to $\phi=A\times A$. Applying the next rule $A\to\mathtt{c}$ yields $\phi=c_{1}\times A$. An index is assigned to the coefficient symbol c, to differentiate multiple coefficients. Finally, applying $A\to\log(x_{1})$ to $\phi=c_{1}\times A$ yields $\phi=c_{1}\log(x_{1})$.
Symbolic Equivalence under a Rewrite System. Rewrite rules are widely used to simplify, rearrange, and reformulate expressions in tasks such as code optimization and automated theorem proving [huet1980equations, DBLP:journals/pacmpl/NandiWZWSASGT21]. A rewrite system provides a principled procedure for transforming expressions by replacing sub-expressions according to predefined patterns.
Formally, a rewrite rule $r_{i}$ is written as $\mathtt{LHS}\leadsto\mathtt{RHS}$”, where the left-hand side ($\mathtt{LHS}$) specifies a pattern to be matched, and the right-hand side ($\mathtt{RHS}$) specifies the substitution applied upon a match. Given a set of rewrite rules $\mathcal{R}=\{r_{1},r_{2},r_{3},\ldots\}$, the symbolic equivalence relation $\equiv_{\mathcal{R}}$ induced by $\mathcal{R}$ is defined as follows: for two symbolic expressions $\phi_{1}$ and $\phi_{2}$,
@@
where $\phi_{1}\Rightarrow^{*}\phi_{2}$ means that $\phi_{1}$ can be transformed into $\phi_{2}$ by applying a finite sequence of rewriting using $\mathcal{R}$. In other words, two expressions are equivalent under $\mathcal{R}$ if one can be transformed into the other via repeated rewriting.
For example, consider the rewrite rule $\log(a\times b)\leadsto\log(a)+\log(b)$ (denoted as $r_{1}$), where $a$ and $b$ are placeholders for arbitrary sub-expressions. Since $\phi_{1}=\log(x_{1}^{3}x_{2}^{2})$ can be transformed into $\phi_{2}=\log(x_{1}^{3})+\log(x_{2}^{2})$ by applying $r_{1}$ once, $\phi_{1}$ and $\phi_{2}$ are symbolically equivalent under $r_{1}$.
In this work, the known mathematical identities listed in Table 3 (in the appendix) are encoded as rewrite rules in $\mathcal{R}$. Section 3.1 applies these rewrite rules over the e-graph to generate a batch of symbolically equivalent expressions via matching (Figure 1b) and substitution (Figure 1c) operations. Implementation details of rewrite rules are provided in Appendix A.2.
Symbolic Regression (SR) posits that experimental data are generated by an underlying closed-form expression, an assumption that is widely adopted across the sciences [doi:10.1126/sciadv.abq0279]. Given a dataset ${D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, the goal is to find an optimal expression $\phi^{*}$ that minimizes the loss:
$\phi^{*}\leftarrow\arg\min_{\phi}\;\frac{1}{N}\sum_{i=1}^{N}\ell(\phi(\mathbf{x}_{i},\mathbf{c}),y_{i}),$
where function $\ell$ measures the discrepancy between the prediction $\phi(\mathbf{x}_{i},\mathbf{c})$ and the ground truth $y_{i}$. The coefficients $\mathbf{c}$ in $\phi$ are typically optimized on the training data D{D} using numerical optimizers such as BFGS [fletcher2000practical]. This problem is NP-hard [journal/tmlr/virgolin2022], posing a major challenge for SR algorithms. Recent efforts to mitigate this challenge are reviewed in Section 4.
@@ 3.1 Egg: Equality graph for Grammar-based Symbolic Expression
Enumerating all equivalent variants of a symbolic expression is combinatorially expensive. Storing these variants explicitly is time-consuming and memory-inefficient. To mitigate this scalability challenge, we adopt the recently proposed Equality graph (i.e, E-graph) data structure [DBLP:journals/pacmpl/WillseyNWFTP21, DBLP:journals/jar/WaldmannTRB22], which compactly represents the set of equivalent expressions by sharing common subexpressions. We extend E-graphs to support grammar-based symbolic expressions, noted as Egg, facilitating unified integration with symbolic regression algorithms.
@@
Definition. An e-graph consists of a collection of equivalence classes, called e-classes. Each e-class contains a set of e-nodes representing symbolically equivalent sub-expressions [DBLP:journals/pacmpl/WillseyNWFTP21]. Each e-node encodes a mathematical operation and references a list of child e-classes corresponding to its operands. Edges always point from an e-node to e-classes.
@@
Figure 1(d) shows an example e-graph. The color-highlighted part is an e-class (in dashed box) containing two e-nodes (in solid boxes): $A\to\log(A)$ and $A\to A+A$. The two e-nodes represent logarithmic operation and addition, respectively. The e-node $A\to\log(A)$ has a single outgoing edge to its child e-class $A\to A\times A$, because $\log(\cdot)$ operator is unary.
Figure 1: Example e-graph construction by applying rewrite rule $\log(a\times b)\leadsto\log(a)+\log(b)$ to an e-graph representing the expression $\log(x_{1}^{3}x_{2}^{2})$. (a) The initialized e-graph consists of e-classes (dashed boxes), each containing equivalent e-nodes (solid boxes). Edges connect e-nodes to their child e-classes. (b) The matching step identifies the e-nodes that match the $\mathtt{LHS}$ of the rule. (c) The substitution step adds new e-classes and edges corresponding to the $\mathtt{RHS}$ to the e-graph. (d) The merging step consolidates equivalent e-classes. See Example 3.2 for a detailed explanation.
@@
Construction. In this work, an e-graph is initialized with a sequence of production rules, representing an input expression. Each rewrite rule (as defined in section 2) is applied by matching its left-hand side ($\mathtt{LHS}$) pattern against the current e-graph, which involves traversing all e-classes to identify subexpressions that match $\mathtt{LHS}$. For every successful match, new e-classes and e-nodes corresponding to the right-hand side ($\mathtt{RHS}$) of the rule are created. A merge operation is applied to incorporate the new e-class with the matched e-class, thereby preserving the structure of known equivalences. This process, known as equality saturation, iteratively applies pattern matching and merging until either no further rules can be applied or a maximum number of iterations is reached.
Example 3.1.
@@
Figure 1 shows an example e-graph construction with the rewrite rule $\log(a\times b)\leadsto\log(a)+\log(b)$, where $a$ and $b$ are placeholders for arbitrary sub-expressions. The e-graph is initialized with an expression $\phi=\log(x_{1}^{3}x_{2}^{2})$ in Figure 1(a). The $\mathtt{LHS}$ is matched against the color-highlighted e-classes in Figure 1(b) with $a=x_{1}^{3}$ and $b=x_{2}^{2}$. The substitution step constructs e-classes and e-nodes that represent $\mathtt{RHS}$, which are color-highlighted in Figure 1(c). The newly created e-class is merged with the matched e-class in Figure 1(d). The resulting e-graph represents two equivalent expressions: $\log(x_{1}^{3}x_{2}^{2})$ and $\log(x_{1}^{3})+\log(x_{2}^{2})$. The e-graph in Figure 1(d) saves memory by storing two sub-expressions $x_{1}^{3}$ and $x_{2}^{2}$ only once. Additional Egg visualizations on more complex expressions are provided in Appendix D.1.
@@
Extraction. After the e-graph is saturated, an extraction step is performed to obtain $K$ representative expressions [DBLP:journals/pacmpl/GoharshadyLP24]. Because an e-graph encodes up to an exponential number of equivalent expressions, exhaustive enumeration is computationally infeasible. We therefore adopt two practical strategies: (1) cost-based extraction, which selects several simplified expressions by minimizing a user-defined cost function over operators and variables [rEGGression], and (2) random-walk sampling, which generates a batch of expressions by stochastically traversing the e-graph. A detailed explanation of extraction is provided in Appendix A.3.2.
Interaction with SR algorithms. Prior research has leveraged e-graphs, based on the principle of Occam’s razor, to obtain the most simplified and least-cost equivalent form [de2025equality]. In this study, however, equivalence-aware learning (detailed in Section 3.2) is encouraged by explicitly exposing SR algorithms to an extra subset of equivalent variants.
The Egg module is primarily used to generate a subset of equivalent variants—i.e., expressions that represent the same mathematical function under a set of math identities. Given a sequence of production rules predicted by an SR algorithm, Egg constructs the e-graph and then performs extraction (via random-walk sampling) to return a batch of equivalent sequences. Egg is implemented for grammar-based symbolic expressions in pure Python, following the original e-graph paper [DBLP:journals/pacmpl/WillseyNWFTP21]. Implementation of Egg is provided in Appendix A.
3.2 Embedding Symbolic Equivalence into Symbolic Regression via Egg